OpenStudy (anonymous):

Solve: 5(x + 6) 3x - 4

OpenStudy (anonymous):

Is that suppose to be an equal sign between: 5(x+6) = 3x - 4?

OpenStudy (anonymous):

5x + 6 = 3x - 4 2x = -10 x = -5

OpenStudy (anonymous):

dont they have a sign in between !?

OpenStudy (jiteshmeghwal9):

Distribute 5 btw x+6 at L.H.S. as 5(x)+5(6)=3x-4 5x+30=3x+4 now transpose variables at one side & constants at one side as 5x-3x+30=3x-3x+4 2x+30=4 2x=4-30 2x=-26 x=-26/2=-13

OpenStudy (anonymous):

so they have a = sign between them,well this is pretty easy then ! 5x+30=3x-4 5x-3x=-4-30 2x=-34 x=-34/2 x=-17